### v1

# Extended Circular Nim

Abstract
Circular nim $CN(m,k)$ is a variant of nim, in which there are $m$ piles of tokens arranged in a circle and each player, in their turn, chooses at most $k$ consecutive piles in the circle and removes an arbitrary number of tokens from each pile. The player must remove at least one token in total. For some cases of $m$ and $k$, closed formulas to determine which player has a winning strategy have been found. Almost all cases are still open problems. In this paper, we consider a variant of circular nim, extended circular nim. In extended circular nim $ECN(m_{S},k)$, there are $m$ piles of tokes arranged in a circle. $S$ is a set of positive integers less than or equal to half of $m$. Each player, in their turn, chooses at most $k$ piles selected every $s$-th pile in a circle for an $s\in S$. We find some closed formulas to determine which player has a winning strategy for the cases where the number of piles is no more than eight.
1 Introduction
In this paper, we consider a generalization of a ruleset, circular nim. In circular nim $CN(m,k)$, there are $m$ piles of tokens arranged in a circle. Each player, in their turn, chooses at most $k$ consecutive piles in the circle and removes an arbitrary number of tokens from each pile. The player must remove at least one token in total. This game has been studied in some articles and winning positions for the previous player are characterized for some cases of $(m,k)$.
In this paper, we consider extended circular nim $ECN(m_{S},k)$, a generalization of circular nim. In $ECN(m_{S},k)$, there are $m$ piles of tokes arranged in a circle. $S$ is a set of positive integers less than or equal to half of $m$. Each player, in their turn, chooses at most $k$ piles selected every $s$-th pile in a circle for an $s\in S$. That is, $ECS(m_{\{1\}},k)$ is the same as $CS(m,k).$
1.1 Combinatorial game theory
In order to study games such as extended circular nim, we need to introduce some background of combinatorial game theory (CGT). Combinatorial game theory studies two-player perfect information games with no chance. In this paper, we consider impartial games, that is, in every position, both players have the same sets of options. We also assume that the games are short, that is, the game must end in a finite number of moves, and under normal play convention, that is, the player who cannot move is the loser.
With these assumptions, the player who has a winning strategy is uniquely determined for each position. A position in which the next player has a winning strategy is called an $\mathcal{N}$-position and a position in which the previous player has a winning strategy is called a $\mathcal{P}$-position.
It is well-known that if all positions of an impartial ruleset are separated into two sets $P$ and $N$, which satisfy the following properties, then $P$ and $N$ are the set of $\mathcal{P}$-positions and $\mathcal{N}$-positions, respectively:
  • All terminal positions are in $P$.
  • For any position in $P$, all the options of the position are in $N$.
  • For any position in $N$, at least one option of the position is in $P$.
We will use this fact for some proof.
1.2 Some impartial rulesets
1.2.1 Nim and Moore’s nim
In early studies, the characterization of $\mathcal{P}$-positions is considered in a simple formula.
As examples of rulesets in which the sets of $\mathcal{P}$-positions are characterized, the following two rulesets, nim and Moore’s nim are well-known.
Definition 1.
In nim, there are some piles of tokens. The player, in their turn, chooses one pile and removes an arbitrary positive number of tokens from the pile.
Note that since we assume that the winner is determined by normal play convention, the player who removes the last token is the winner.
A position in nim is denoted as $(n_{0},n_{1},\ldots,n_{m-1})$ when there are $m$ piles and each pile has $n_{0},n_{1},\ldots,n_{m-1}$ tokens. Let $\oplus$ be the exclusive OR operator for binary notation.
By using the following theorem, we can easily determine which player has a winning strategy in nim.
Theorem 1 ([1]).
A position $(n_{0},n_{1},\ldots,n_{m-1})$ in nim is a $\mathcal{P}$-position if and only if $n_{0}\oplus n_{1}\oplus\cdots\oplus n_{m-1}=0$.
Moore’s nim is a generalization of nim.
Definition 2.
In Moore’s nim ${\rm MN}(m,k)$, there are $m$ piles of tokens. Each player, in their turn, chooses at most $k$ piles and removes an arbitrary number of tokens from each pile. The player must remove at least one token in total.
When $k=1,$ Moore’s nim is the original nim. Thus, Moore’s nim is a generalization of nim. The $\mathcal{P}$-positions of this ruleset are characterized by the following theorem.
Theorem 2 ([8]).
A position $(n_{0},n_{1},\ldots,n_{m-1})$ in Moore’s nim ${\rm MN}(m,k)$, is a $\mathcal{P}$-position if and only if when each $n_{i}$ is expanded in the binary form, the summands of each digit contain a multiple of (k + 1) number of 1s.
This theorem is a generalization of Theorem 1.
Note that when $k=m-1$, a position $(n_{0},n_{1},\ldots,n_{m-1})$ in Moore’s nim ${\rm MN}(m,m-1)$ is a $\mathcal{P}$-position if and only if $n_{0}=n_{1}=\cdots=n_{m-1}$.
1.2.2 Simplicial nim
Simplicial nim is a generalization of nim and Moore’s nim.
Let $V$ be a finite set of vertices. A simplicial complex $\Delta$ on $V$ is a subset of $2^{\mid V\mid}$ which satisfies the following:
  • For any element $v\in V,$$\{v\}\in\Delta$
  • If $F\in\Delta$ and $G\subseteq F,$ then $G\in\Delta$.
An element of $\Delta$ is termed a face of $\Delta$. Ehrenborg and Steingrímsson studied nim on simplicial complexes or simplicial nim [4]. Let $\Delta$ be a simplicial complex of a finite set $V$. Each vertex has several tokens. In nim on $\Delta$, the players, in turn, choose a non-empty face $F$ in $\Delta$ and arbitrarily remove any positive number of tokens from all vertices in $F$. The original nim and moore’s nim can be considered as special cases of this game. Ehrenborg and Steingrímsson did not characterize the $\mathcal{P}$-positions of nim on $\Delta$ without restrictions but discovered good constructions and characterized $\mathcal{P}$-positions for certain cases. For example, they proved that circuits, which are subsets of the vertex set and not faces of the simplicial complex, but all proper subsets of them are faces, can be used for characterizing $\mathcal{P}$-positions. Horrocks [6] and Penn [10] also researched this topic.
For nonnegative integers $n_{0},n_{1},\ldots,n_{m-1},$ a position $(n_{0},n_{1},\ldots,n_{m-1})$ of nim on $\Delta$ is a position such that there are $n_{i}$ tokens on $v_{i}$ for each $i$.
Example 1.
Let $V=(v_{0},v_{1},v_{2})$ and $\Delta=\{\{v_{0}\},\{v_{1}\},\{v_{2}\},\{v_{0},v_{1}\}\}.$ Consider the position $A=(3,4,5)$ of nim on $\Delta$. In this case, the position $B=(1,2,5)$ is an option of $A$ because $\{v_{0},v_{1}\}$ is a face of $\Delta$ but the position $C=(3,3,3)$ is not an option of $A$ because $\{v_{1},v_{2}\}$ is not a face of $\Delta$.
1.2.3 Circular nim
Now we can define circular nim as a special case of simplicial nim.
Definition 3.
Let $m$ and $k$ be positive integers with $k\leq m$. Circular nim ${\rm CN}(m,k)$ is nim on a simplicial complex $\Delta$, where $V=\{v_{0},\ldots,v_{m-1}\}$ and $\Delta=\bigcup_{i=0}^{m-1}\bigcup_{j=i}^{i+k-1}\{\{v_{i},v_{((i+1)\bmod m)},\ldots,v_{(j\bmod m)}\}\}$.
That is, in circula nim, there are $m$ piles of tokens arranged in a circle and each player, in their turn, chooses at most $k$ consecutive piles in the circle and removes an arbitrary number of tokens from each pile, where in total at least one token must be removed [^2] Note that in circular nim and extended circular nim defined in the next section, a pile is considered to be still existing when counting $k$ consecutive piles even if all tokens on the pile are removed. In contrast, in shrinking circular nim, a variant of circular nim discussed in [9], an empty pile is ignored.. When $k=1$, ${\rm CN}(m,k)$ is the same ruleset as $m$-pile nim. Also, when $k=m-1$, ${\rm CN}(m,k)$ is the same ruleset as Moore’s nim ${\rm MN}(m,m-1)$. Furthermore, when $k=m$, the player can select all of the piles; thus we can consider this ruleset as a single-pile nim. Therefore, these cases are trivial and other cases are studied in [4, 6, 2, 3] as follows. For any $M=(n_{0},\ldots,n_{m-1})$, we say $M\in_{\circlearrowleft}P$ if there exists $i(<m)$ such that $(n_{i},n_{((i+1)\bmod m)},\ldots,n_{((i+m-1)\bmod m)})\in P$ or $(n_{i},n_{((i+m-1)\bmod m)},\ldots,n_{((i+1)\bmod m)})\in P$.
  • Consider a position $M=(n_{0},n_{1},n_{2},n_{3})$ in ${\rm CN}(4,2)$. $M$ is a $\mathcal{P}$-position if and only if $n_{0}=n_{2}$ and $n_{1}=n_{3}$.
  • Consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4})$ in ${\rm CN}(5,2)$. Then $M$ is a $\mathcal{P}$-position if and only if $M\in_{\circlearrowleft}P=\{(n_{0},n_{1},n_{2},n_{3},n_{4})\mid n_{0}=\max(M),n_{0}+n_{1}=n_{2}+n_{3},n_{1}=n_{4}\}$.
  • Consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4})$ in ${\rm CN}(5,3)$. Then $M$ is a $\mathcal{P}$-position if and only if $M\in_{\circlearrowleft}P=\{(n_{0},n_{1},n_{2},n_{3},n_{4})\mid n_{0}=0,n_{1}=n_{2}+n_{3}=n_{4}\}.$
  • Consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})$ in ${\rm CN}(6,3)$. Then $M$ is a $\mathcal{P}$-position if and only if $M\in_{\circlearrowleft}P=\{(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})\mid n_{0}+n_{1}=n_{3}+n_{4},n_{1}+n_{2}=n_{4}+n_{5}\}$.
  • Consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})$ in ${\rm CN}(6,4)$. Then, $M$ is a $\mathcal{P}$-position if and only if $M\in_{\circlearrowleft}P=\{(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})\mid n_{0}=\min(M),n_{0}+n_{1}=n_{3}+n_{4},n_{1}+n_{2}=n_{4}+n_{5},n_{0}\oplus n_{2}\oplus n_{4}=~{}0\}$.
  • Consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6})$ in ${\rm CN}(7,4)$. Then, $M$ is a $\mathcal{P}$-position if and only if $M\in_{\circlearrowleft}P=P_{1}\cup P_{2}\cup P_{3}\cup P_{4}$, where
    $P_{1}$ $=$ $\{(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6})\mid n_{0}=n_{1}=0,n_{3}+n_{4}+n_{5}=n_{2}=n_{6}>0\},$
    $P_{2}$ $=$ $\{(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6})\mid n_{0}=n_{1}=n_{2}=n_{3}=n_{4}=n_{5}=n_{6}\},$
    $P_{3}$ $=$ $\{(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6})\mid n_{0}=n_{1},n_{2}=n_{6},n_{3}=n_{5},n_{0}+n_{2}=n_{3}+n_{4},$
    $0<n_{0}<n_{4},n_{0}=\min(M)\},$
    $P_{4}$ $=$ $\{(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6})\mid n_{0}=n_{5},n_{1}+n_{2}=n_{3}+n_{4}=n_{6}+n_{0},$
    $n_{0}=\min(M),n_{0}<\min(\{n_{1},n_{4}\}),n_{0}<\max(\{n_{2},n_{3}\})\}.$
  • Consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6},n_{7})$ in ${\rm CN}(8,6)$. Then, $M$ is a $\mathcal{P}$-position if and only if $M\in_{\circlearrowleft}P=\{(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6},n_{7})\mid n_{0}=0,n_{1}=n_{2}+n_{3}=n_{5}+n_{6}=n_{7},n_{4}=\min(\{n_{1},n_{2}+n_{6}\})\}$
For the other cases of circuar nim, it is still open problems to find a closed formula for determining the set of $\mathcal{P}$-positions.
As shown here, in circular nim, the formulas for $\mathcal{P}$-positions are completely different despite the rules are similar. This is the reason why almost all cases are still unsolved problems.
In this paper, we aim to reexamine circular nim from a broader perspective by extending its rules and explore the relationship between the rules and the formula that determines $\mathcal{P}$-positions.
1.3 Sums of positions
We also use the following two types of sums of positions.
1.3.1 Disjunctive sum
Definition 4.
Let $G$ and $H$ be positions of impartial rulesets. The disjunctive sum of $G$ and $H$ , which is denoted as $G+H$, is a position such that the set of its all options is
$\{G^{\prime}+H,G+H^{\prime}\mid G^{\prime}\text{ is an option of }G,H^{\prime}\text{ is an option of }H\}.$
That is, in the disjunctive sum of positions, the current player chooses exactly one component and makes a legal move on the selected component. One can determine which player has a winning strategy by using Sprague-Grundy values.
Definition 5.
[…]
$\mathcal{G}(G)={\rm mex}(\{\mathcal{G}(G^{\prime})\mid G^{\prime}\text{is an option of }G\}).$
Here, ${\rm mex}(S)=\mathbb{N}_{0}\setminus S$ and $\mathbb{N}_{0}$ is the set of all nonnegative integers.
Theorem 3 ([12, 5]).
The followings hold:
  • Let $G$ be a position of an impartial ruleset. Then $G$ is a $\mathcal{P}$-position if and only if $\mathcal{G}(G)=0$.
  • Let $G$ and $H$ be positions of impartial rulesets. Then, $\mathcal{G}(G+H)=\mathcal{G}(G)\oplus\mathcal{G}(H).$
1.3.2 Selective sum
Definition 6.
Let $G$ and $H$ be positions of impartial rulesets. The selective sum of $G$ and $H$ , which is denoted as $G\vee H$, is a position such that the set of its all options is
$\{G^{\prime}\vee H,G\vee H^{\prime},G^{\prime}\vee H^{\prime}\mid G^{\prime}\text{ is an option of }G,H^{\prime}\text{ is an option of }H\}.$
That is, in the selective sum of positions, the current player chooses at least one component and makes a legal move on each selected component.
Theorem 4 ([11]).
Let $G$ and $H$ be positions of impartial rulesets. $G\vee H$ is a $\mathcal{P}$-position if and only if both $G$ and $H$ are $\mathcal{P}$-positions.
2 Main results
2.1 Extended circular nim
Extended circular nim $ECN(m_{S},k)$ is defined as follows.
Definition 7.
Let $m$ and $k$ be positive integers with $k\leq m$ and let $S$ be a set of positive integers that are not more than half of $m$. Extended circular nim ${\rm ECN}(m_{S},k)$ is nim on a simplicial complex $\Delta$, where $V=\{v_{0},\ldots,v_{m-1}\}$ and $\Delta=\bigcup_{s\in S}\bigcup_{i=0}^{m-1}\bigcup_{j=0}^{k-1}\{\{v_{i},v_{((i+s)\bmod m)},\ldots,v_{((i+js)\bmod m)}\}\}$.
That is, in extended circular nim $ECN(m_{S},k)$, there are $m$ piles of tokes arranged in a circle. $S$ is a set of positive integers less than or equal to half of $m$. Each player, in their turn, chooses at most $k$ piles selected every $s$-th pile in a circle for an $s\in S$.
Figure 1 shows piles of a position in ${\rm ECN}(6_{\{1,2\}},3)$. Selecting piles $A,B,$ and $C$ is legal because the player selects every first pile (connected black lines), also, selecting piles $A,C$, and $E$ is legal because the player selects every second pile (connected green lines). However selecting piles $A,B$ , and $D$ is illegal in this ruleset.
Figure 1: ${\rm ECN}(6_{1,2},3)$
Definition 8.
Let $\Gamma_{1}$ and $\Gamma_{2}$ be rulesets. $\Gamma_{1}$ and $\Gamma_{2}$ are isomorphic if there is a bijection map $f$ from the set of all positions in $\Gamma_{1}$ to the set of all positions in $\Gamma_{2}$ such that if $g_{1},g_{2}$ are positions in $\Gamma_{1}$ and $g_{2}$ is an option of $g_{1},$ then $f(g_{2})$ is an option of $f(g_{1})$ and vice versa.
Note that from the definition of circular nim and extended circular nim, ${\rm ECN}(m_{\{1\}},k)$ and ${\rm CN}(m,k)$ are isomorphic.
Theorem 5.
Let $a$ be an integer which satisfies $a\geq 4$.
When an integer $b\in S$ is prime to $a$, then ${\rm ECN}(a_{S},a-1)$ is isomorphic to ${\rm MN}(a,a-1)$. Otherwise, it is isomorphic to ${\rm ECN}(a_{S},a-2)$.
When an integer $b\in S$ is prime to $a$, then ${\rm ECN}(a_{S},a)$ is isomorphic to ${\rm MN}(a,a)$; that is, this can be regarded as a one-pile nim. Otherwise, it is isomorphic to ${\rm ECN}(a_{S},a-2)$.
Proof.
Consider ${\rm ECN}(a_{S},a-1)$. If $b\in S$ is prime to $a$, the player can choose arbitrary $a-1$ piles. Therefore, ${\rm ECN}(a_{S},a-1)$ is isomorphic to ${\rm MN}(a,a-1)$. Otherwise, the player can choose at most $\frac{a}{2}$ piles. Since $\frac{a}{2}\leq a-2$, this ruleset is isomorphic to ${\rm ECN}(a_{S},a-2)$.
Similarly, ${\rm ECN}(a_{S},a)$ is isomorphic to ${\rm MN}(a,a)$ when an integer $b\in S$ is pritme to $a$ and otherwise ${\rm ECN}(a_{S},a)$ is isomorphic to ${\rm ECN}(a_{S},a-2)$. ∎
Table 1 summarizes these discussions. From them, when considering extended circular nim, we do not need to consider ${\rm ECN}(a_{S},1),{\rm ECN}(a_{S},a-1),{\rm ECN}(a_{S},a)$ independently. In this paper, we consider ${\rm ECN}(a_{S},i)(1<i<a-1)$.
Ruleset Result
${\rm ECN}(a_{S},1)$ Isomorphic to $a$-pile nim
${\rm ECN}(a_{S},a-1)$ Theorem 5
${\rm ECN}(a_{S},a)$ Theorem 5
Table 1: Results for easy cases
2.2 Extended circular nim with less than six piles
We can characterize $\mathcal{P}$-positions in extended circular nim with less than six piles by using early results.
Consider ${\rm ECN}(4_{S},2)$.
When $S=\{1\}$, the ruleset is already considered as circular nim.
When $S=\{2\},$ we can combine each diagonal piles as single piles. Therefore, the position can be considered as a position in two-pile nim.
Consier ${\rm ECN}(5_{S},k)(k=2,3)$. When $S=\{1\}$, the ruleset is already considered as circular nim. As shown in Fig. 2, ${\rm ECN}(5_{\{2\}},i)$ is isomorphic to ${\rm ECN}(5_{\{1\}},i)$. When $S=\{1,2\}$, arbitrary two piles can be selected in ${\rm ECN}(5_{\{1,2\}},2)$ and arbitrary three piles can be selected in ${\rm ECN}(5_{\{1,2\}},3)$. Therefore, these rulesets are isomorphic to ${\rm MN}(5,2)$ and ${\rm MN}(5,3)$, respectively.
Table 2 summarizes these resules.
Figure 2: ${\rm ECN}(5_{\{1\}},i)$ and ${\rm ECN}(5_{\{2\}},i)$ are isomorphic
Ruleset Result
${\rm ECN}(4_{1},2)$ Shown in [2]
${\rm ECN}(4_{2},2)$ Regarded as $2$-pile nim
${\rm ECN}(4_{1,2},2)$ Isomorphic to ${\rm MN}(4,2)$
${\rm ECN}(5_{1},2)$ Shown in [4]
${\rm ECN}(5_{1},3)$ Shown in [4]
${\rm ECN}(5_{2},i)(2\leq i\leq 3)$ Isomorphic to ${\rm ECN}(5_{1},i)$
${\rm ECN}(5_{1,2},i)(2\leq i\leq 3)$ Isomorphic to ${\rm MN}(5,i)$
Table 2: Results for extended circular nim with less than six piles
2.3 Extended circular nim with six piles
When there are six or more piles, some cases are not isomorphic to rulesets which are already studied. We show the characterizations of $\mathcal{P}$-positions for such rulesets.
First, consider ${\rm ECN}(6_{\{2\}},2)$. This ruleset can be considered as a disjunctive sum of two positions in ${\rm MN}(3,2)$. In general, closed formula for Sprague-Grundy values of Moore’s nim is not found, but for the case of ${\rm MN}(n,n-1)$, the formula for Sprague-Grundy values is found in [7]. Thus, in ${\rm ECN}(6_{\{2\}},2)$, the set of $\mathcal{P}$-positions can be characterized.
Next, we consider ${\rm ECN}(6_{\{1,2\}},2)$.
Theorem 6.
Consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})$ in ${\rm ECN}(6_{\{1,2\}},2)$. $M$ is a $\mathcal{P}$-position if and only if $M\in P=\{(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})\mid n_{0}\oplus n_{3}=n_{1}\oplus n_{4}=n_{2}\oplus n_{5}\}$.
Proof.
Let $N$ be the set of positions which are not in $P$. The terminal position is in $P$.
Next, consider $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})\in P$. Let $k=n_{0}\oplus n_{3}=n_{1}\oplus n_{4}=n_{2}\oplus n_{5}$. Then, for any $M^{\prime}=(n^{\prime}_{0},n^{\prime}_{1},n^{\prime}_{2},n^{\prime}_{3},n^{\prime}_{4},n^{\prime}_{5})$ which is an option of $M$, let $a_{1}=n^{\prime}_{0}\oplus n^{\prime}_{3},a_{2}=n^{\prime}_{1}\oplus n^{\prime}_{4},a_{3}=n_{2}^{\prime}\oplus n^{\prime}_{5}$. At least one $a_{i}$ , $a_{i}\neq k$ and there exists an $a_{j}$, $a_{j}=k$. Therefore, $M^{\prime}\not\in P$.
Lastly, consider $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})\in N$. Let $k=\min(n_{0}\oplus n_{3},n_{1}\oplus n_{4},n_{2}\oplus n_{5})$. Without loss of generality, $k=n_{0}\oplus n_{3}$. Then, at least one of $n_{1}\oplus n_{4}>k$ and $n_{2}\oplus n_{5}>k$ holds. In addition, each of $(n_{1}),(n_{4}),(n_{2}),(n_{5}),(n_{1},n_{2}),(n_{1},n_{5}),(n_{4},n_{2})$ and $(n_{4},n_{5})$ can be reduced in one move in ${\rm ECN}(6_{\{1,2\}},2)$. Therefore, there is an option $M^{\prime}$ of $M$ which satisfies $M^{\prime}=(n_{0},n^{\prime}_{1},n^{\prime}_{2},n_{3},n^{\prime}_{4},n^{\prime}_{5})$ and $n_{0}\oplus n_{3}=n^{\prime}_{1}\oplus n^{\prime}_{4}=n^{\prime}_{2}\oplus n^{\prime}_{5}$. ∎
Theorem 7.
Consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})$ in ${\rm ECN}(6_{\{1,2\}},3)$. Then, $M$ is a $\mathcal{P}$-position if and only if $M\in P=\{(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})\mid n_{0}=n_{3},n_{1}=n_{4},n_{2}=n_{5}\}$.
Proof.
Consider a position $(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})$ in ${\rm ECN}(6_{\{1,2\}},3)$. This position is a selective sum of $(n_{0},n_{3})$, $(n_{1},n_{4})$, and $(n_{2},n_{5})$, which are positions of two-pile nim. Therefore, $n_{0}\oplus n_{3}=n_{1}\oplus n_{4}=n_{2}\oplus n_{5}=0$, that is, $n_{0}=n_{3},n_{1}=n_{4},n_{2}=n_{5}$ is a necessary and sufficient condition for $(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})$ is a $\mathcal{P}$-position.
Theorem 8.
Let $P$ be the set of positions which satisfies the following:
  • Assume that $(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})\in P$. When each $n_{i}$ is expanded in the binary form, the summands of each digit contain four or no $1$s. In addition, for each digit contain four $1$s, the $1$s are in $(n_{0},n_{1},n_{3},n_{4})$ or $(n_{0},n_{1},n_{2},n_{4})$.
Consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})$ in ${\rm ECN}(6_{\{1,2\}},4)$. Then, M is a $\mathcal{P}$-position if and only if $M\in_{\circlearrowleft}P$.
Proof.
This can be proven by using an early result on simplicial complex. We can use Theroem 4.4 in [4].
Theorem 9.
Consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})$ in ${\rm ECN}(6_{\{1,3\}},2)$. Then, $M$ is a $\mathcal{P}$-position if and only if $M\in P=\{(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})\mid n_{0}\oplus n_{2}\oplus n_{4}=n_{1}\oplus n_{3}\oplus n_{5}=0\}$.
Proof.
Consider a position $(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})$ in ${\rm ECN}(6_{\{1,3\}},2)$. This position is a selective sum of $(n_{0},n_{2},n_{4})$ and $(n_{1},n_{3},n_{5})$, which are positions of three-pile nim. Therefore, $(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})$ is a $\mathcal{P}$-position if and only if $n_{0}\oplus n_{2}\oplus n_{4}=n_{1}\oplus n_{3}\oplus n_{5}=0$ ∎
Theorem 10.
Consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})$ in $ECN(6_{\{2,3\}},3)$. $M$ is a $\mathcal{P}$-position if and only if $M\in_{\circlearrowleft}P=\{(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})\mid n_{0}+n_{2}+n_{4}=n_{1}+n_{3}+n_{5},n_{0}\oplus n_{1}\oplus n_{2}=0,n_{0}\leq n_{3},n_{1}\leq n_{4},n_{2}\leq n_{5}\}$.
Proof.
Let $N=\{M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})\mid M\not\in_{\circlearrowleft}P\}$. The terminal position is in $P$.
Next, consider $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})\in P$. Then, $n_{0}\leq n_{3},n_{1}\leq n_{4},$ and $n_{2}\leq n_{5}$ hold. Let $M^{\prime}=(n^{\prime}_{0},n^{\prime}_{1},n^{\prime}_{2},n^{\prime}_{3},n^{\prime}_{4},n^{\prime}_{5})$ be an option of $M$. When the player reduces some of $n_{0},n_{2},n_{4}$, or reduces some of $n_{1},n_{3},n_{5}$, $n^{\prime}_{0}+n^{\prime}_{2}+n^{\prime}_{4}\neq n^{\prime}_{1}+n^{\prime}_{3}+n^{\prime}_{5}$ holds and $M^{\prime}\not\in P$. For the case that the player reduces the same number from $n_{0}$ and $n_{3}$, $n^{\prime}_{0}+n^{\prime}_{2}+n^{\prime}_{4}=n^{\prime}_{1}+n^{\prime}_{3}+n^{\prime}_{5}$ is satisfied but $n^{\prime}_{0}\oplus n^{\prime}_{1}\oplus n^{\prime}_{2}=n^{\prime}_{0}\oplus n_{1}\oplus n_{2}\neq 0$ holds. We show $M^{\prime}\not\in_{\circlearrowleft}P$ by contradiction:
  • (i)
    Assume that $n^{\prime}_{1}\oplus n^{\prime}_{2}\oplus n^{\prime}_{3}=0$. If $n_{0}=n_{3}$, then $n^{\prime}_{0}=n^{\prime}_{3}$, which contradicts to $n^{\prime}_{0}\oplus n^{\prime}_{1}\oplus n^{\prime}_{2}\neq 0$. In addition, $n_{0}<n_{3}$ does not satisfy $n^{\prime}_{3}\leq n^{\prime}_{0}$.
  • (ii)
    Assume that $n^{\prime}_{2}\oplus n^{\prime}_{3}\oplus n^{\prime}_{4}=n_{2}\oplus n^{\prime}_{3}\oplus n_{4}=0$. In order to satisfy $M^{\prime}\in P$, it is necessary to be $n^{\prime}_{3}\leq n^{\prime}_{0},n_{4}\leq n_{1}$. From the assumption, $n^{\prime}_{0}\leq n^{\prime}_{3},n_{1}\leq n_{4}$, so $n^{\prime}_{0}=n^{\prime}_{3},n_{1}=n_{4}$ hold. However, then we have $n_{2}\oplus n^{\prime}_{3}\oplus n_{4}=n_{2}\oplus n^{\prime}_{0}\oplus n_{1}=0$, which contradicts to $n^{\prime}_{0}\oplus n_{1}\oplus n_{2}\neq 0$.
  • (iii)
    Assume that $n^{\prime}_{3}\oplus n^{\prime}_{4}\oplus n^{\prime}_{5}=n^{\prime}_{3}\oplus n_{4}\oplus n_{5}=0$. In order to satisfy $M^{\prime}\in P$, it is necessary to be $n^{\prime}_{3}\leq n^{\prime}_{0},n_{4}\leq n_{1},n_{5}\leq n_{2}$. From the assumption, $n^{\prime}_{0}\leq n^{\prime}_{3},n_{1}\leq n_{4},n_{2}\leq n_{5}$, so $n^{\prime}_{0}=n^{\prime}_{3},n_{1}=n_{4},n_{2}=n_{5}$ hold. This contradicts to $n^{\prime}_{0}\oplus n_{1}\oplus n_{2}\neq 0$.
Even if we assume that $n^{\prime}_{4}\oplus n^{\prime}_{5}\oplus n^{\prime}_{0}=0$ or $n^{\prime}_{5}\oplus n^{\prime}_{0}\oplus n^{\prime}_{1}=0$, we have a contradiction in a similar way. Therefore, $M^{\prime}\in N$.
For the cases that the same number is reduced from $n_{1}$ and $n_{4}$, or the same number is reduced from $n_{2}$ and $n_{5}$, we have $M^{\prime}\in N$ in similar ways.
Also, for any position $M\not\in P$ but $M\in_{\circlearrowleft}P$, we can show that every option $M^{\prime}\in N$ in a similar way.
Finally, consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5})\in N$.
  • (i)
    Assume that $n_{0}+n_{2}+n_{4}\neq n_{1}+n_{3}+n_{5}$. Without loss of generality, $n_{0}+n_{2}+n_{4}>n_{1}+n_{3}+n_{5}$. Then, $n_{0}>n_{3}$, $n_{2}>n_{5}$, or $n_{4}>n_{1}$ holds. Without loss of generality, we assume that $n_{0}>n_{3}$.
    • (i-a)
      Assume that $n_{2}\leq n_{5},n_{4}\leq n_{1}$.
      • (i-a-1)
        If $n_{2}\oplus n_{4}\leq n_{3}$, then let $n^{\prime}_{3}=n_{2}\oplus n_{4},n^{\prime}_{0}=n_{1}+n^{\prime}_{3}+n_{5}-n_{2}-n_{4}$ and a move changes $(n_{0},n_{3})$ to $(n^{\prime}_{0},n^{\prime}_{3})$ is a legal move, which satisfies $n^{\prime}_{0}+n_{2}+n_{4}=n_{1}+n^{\prime}_{3}+n_{5},n_{2}\oplus n^{\prime}_{3}\oplus n_{4}=0,n_{2}\leq n_{5},n^{\prime}_{3}\leq n^{\prime}_{0},n_{4}\leq n_{1}$.
      • (i-a-2)
        If $n_{2}\oplus n_{4}>n_{3}$, then $n_{2}\geq n_{3}\oplus n_{4}$ or $n_{4}\geq n_{2}\oplus n_{3}$ holds. Without loss of generality, we assume that $n_{2}\geq n_{3}\oplus n_{4}$. Let $n^{\prime}_{2}=n_{3}\oplus n_{4}\leq n_{2}$. If $n_{0}+n^{\prime}_{2}+n_{4}<n_{1}+n_{3}+n_{5}$, then let $n^{\prime}_{5}=n_{0}+n^{\prime}_{2}+n_{4}-n_{1}-n_{3}$. It satisfies $n^{\prime}_{2}\leq n^{\prime}_{5}<n_{5}$ and $n_{0}+n^{\prime}_{2}+n_{4}=n_{1}+n_{3}+n^{\prime}_{5},n^{\prime}_{2}\oplus n_{3}\oplus n_{4}=0,n^{\prime}_{2}\leq n^{\prime}_{5},n_{3}<n_{0},n_{4}\leq n_{1}$ hold. Thus, $M^{\prime}=(n_{0},n_{1},n^{\prime}_{2},n_{3},n_{4},n^{\prime}_{5})$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$. On the other hand, if $n_{0}+n^{\prime}_{2}+n_{4}\geq n_{1}+n_{3}+n_{5}$, then $n^{\prime}_{2}\leq n_{5},n_{4}\leq~{}n_{1}$, so by letting $n^{\prime}_{0}=n_{1}+n_{3}+n_{5}-n^{\prime}_{2}-n_{4}$, we have $n^{\prime}_{0}+n^{\prime}_{2}+n_{4}=n_{1}+n_{3}+n_{5},n^{\prime}_{2}\oplus n_{3}\oplus n_{4}=0,\allowbreak n^{\prime}_{2}\leq n_{5},n_{3}\leq n^{\prime}_{0},n_{4}\leq n_{1}$. Thus, $M^{\prime}=(n^{\prime}_{0},n_{1},n^{\prime}_{2},n_{3},n_{4},n_{5})$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$.
    • (i-b)
      Assume that $n_{2}>n_{5},n_{4}\leq n_{1}$. Note that at least one of $n_{4}\oplus n_{5}\leq n_{3},n_{4}\oplus n_{3}\leq n_{5}$ and $n_{5}\oplus n_{3}\leq n_{4}$ holds. We separate into some cases.
      • (i-b-1)
        Consider the case that $n_{4}\oplus n_{5}\leq n_{3}$ and $n_{2}+n_{4}\leq n_{1}+n_{5}$. Let $n^{\prime}_{3}=n_{4}\oplus n_{5},n^{\prime}_{0}=n_{1}+n_{5}-n_{2}-n_{4}+n^{\prime}_{3}$ , then $n^{\prime}_{0}+n_{2}+n_{4}=n_{1}+n^{\prime}_{3}+n_{5},n^{\prime}_{3}\oplus n_{4}\oplus n_{5}=0,n^{\prime}_{0}\geq n^{\prime}_{3},n_{2}>n_{5},n_{1}\geq n_{4}$ and $M^{\prime}=(n^{\prime}_{0},n_{1},n_{2},n^{\prime}_{3},n_{4},n_{5})\in_{\circlearrowleft}P$. In addition, since $n_{0}+n_{2}+n_{4}>n_{1}+n_{3}+n_{5}$, we have $n_{0}>n_{1}+n_{3}+n_{5}-n_{2}-n_{4}\geq n_{1}+n_{5}+n^{\prime}_{3}-n_{2}-n_{4}=n^{\prime}_{0}$ and $n_{3}\geq n^{\prime}_{3}$. Thus, $M^{\prime}$ is an option of $M$.
      • (i-b-2)
        Consider the case that $n_{4}\oplus n_{5}\leq n_{3}$ and $n_{2}+n_{4}>n_{1}+n_{5}$. Then, if $n_{1}+n_{3}+n_{5}>(n_{4}\oplus n_{5})+n_{2}+n_{4}$, that is, $n_{3}>n_{2}+n_{4}-n_{1}-n_{5}+(n_{4}\oplus n_{5})>(n_{4}\oplus n_{5})$, let $n^{\prime}_{3}=n_{2}+n_{4}-n_{1}-n_{5}+(n_{4}\oplus n_{5}),n^{\prime}_{0}=n_{4}\oplus n_{5}$. We have $n^{\prime}_{0}+n_{2}+n_{4}=n_{1}+n^{\prime}_{3}+n_{5},n_{4}\oplus n_{5}\oplus n^{\prime}_{0}=0,n_{4}\leq n_{1},n_{5}<n_{2},n^{\prime}_{0}<n^{\prime}_{3}$. Thus, $M^{\prime}=(n^{\prime}_{0},n_{1},n_{2},n^{\prime}_{3},n_{4},n_{5})$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$. On the other hand, if $n_{1}+n_{3}+n_{5}\leq(n_{4}\oplus n_{5})+n_{2}+n_{4}$, then since $n_{3}\geq n_{4}\oplus n_{5}$ and $n_{1}\geq n_{4}$, there exists $n^{\prime}_{2}$ which satisfies $n_{5}\leq n^{\prime}_{2}\leq n_{2}$ and $n_{1}+n_{3}+n_{5}=(n_{4}\oplus n_{5})+n^{\prime}_{2}+n_{4}$. Therefore, let $n^{\prime}_{0}=n_{4}\oplus n_{5}$ and we have $n^{\prime}_{0}+n^{\prime}_{2}+n_{4}=n_{1}+n_{3}+n_{5},n_{4}\oplus n_{5}\oplus n^{\prime}_{0}=0,n_{4}\leq n_{1},n_{5}\leq n^{\prime}_{2},n^{\prime}_{0}\leq n_{3}$. Thus, $M^{\prime}=(n^{\prime}_{0},n_{1},n^{\prime}_{2},n_{3},n_{4},n_{5})$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$.
      • (i-b-3)
        Consider the case that $n_{4}\oplus n_{3}\leq n_{5}$ and $n_{0}+n_{4}\leq n_{1}+n_{3}$. Let $n^{\prime}_{5}=n_{4}\oplus n_{3},n^{\prime}_{2}=n_{1}+n_{3}-n_{0}-n_{4}+n^{\prime}_{5}$, then $n_{0}+n^{\prime}_{2}+n_{4}=n_{1}+n_{3}+n^{\prime}_{5},n_{3}\oplus n_{4}\oplus n^{\prime}_{5}=0,n_{0}\geq n_{3},n^{\prime}_{2}\geq n^{\prime}_{5},n_{1}\geq n_{4}$ and $M^{\prime}=(n_{0},n_{1},n^{\prime}_{2},n_{3},n_{4},n^{\prime}_{5})\in_{\circlearrowleft}P$. In addition, since $n_{0}+n_{2}+n_{4}>n_{1}+n_{3}+n_{5}$, we have $n_{2}>n_{1}+n_{3}+n_{5}-n_{0}-n_{4}\geq n_{1}+n^{\prime}_{5}+n_{3}-n_{0}-n_{4}=n^{\prime}_{2}$ and $n_{5}\geq n^{\prime}_{5}$. Thus, $M^{\prime}$ is an option of $M$.
      • (i-b-4)
        Consider the case that $n_{4}\oplus n_{3}\leq n_{5}$ and $n_{0}+n_{4}>n_{1}+n_{3}$. Then, if $n_{1}+n_{3}+n_{5}>(n_{4}\oplus n_{3})+n_{0}+n_{4}$, that is, $n_{5}>n_{0}+n_{4}-n_{1}-n_{3}+(n_{4}\oplus n_{3})>(n_{4}\oplus n_{3})$, let $n^{\prime}_{5}=n_{0}+n_{4}-n_{1}-n_{3}+(n_{4}\oplus n_{3}),n^{\prime}_{2}=n_{4}\oplus n_{3}$. We have $n_{0}+n^{\prime}_{2}+n_{4}=n_{1}+n_{3}+n^{\prime}_{5},n^{\prime}_{2}\oplus n_{3}\oplus n_{4}=0,n^{\prime}_{2}<n^{\prime}_{5},n_{3}<n_{0},n_{4}\leq n_{1}$. Thus, $M^{\prime}=(n_{0},n_{1},n^{\prime}_{2},n_{3},n_{4},n^{\prime}_{5})$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$. On the other hand, if $n_{1}+n_{3}+n_{5}\leq(n_{4}\oplus n_{3})+n_{0}+n_{4}$, then since $n_{5}\geq n_{4}\oplus n_{3}$ and $n_{1}\geq n_{4}$, there exists $n^{\prime}_{0}$ which satisfies $n_{3}\leq n^{\prime}_{0}\leq n_{0}$ and $n_{1}+n_{3}+n_{5}=(n_{4}\oplus n_{3})+n^{\prime}_{0}+n_{4}$. Therefore, let $n^{\prime}_{2}=n_{4}\oplus n_{3}$ and we have $n^{\prime}_{0}+n^{\prime}_{2}+n_{4}=n_{1}+n_{3}+n_{5},n^{\prime}_{2}\oplus n_{3}\oplus n_{4}=0,n^{\prime}_{2}\leq n_{5},n_{3}\leq n^{\prime}_{0},n_{4}\leq n_{1}$. Thus, $M^{\prime}=(n^{\prime}_{0},n_{1},n^{\prime}_{2},n_{3},n_{4},n_{5})$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$.
      • (i-b-5)
        Consider the case that $n_{5}\oplus n_{3}\leq n_{4}$. Let $n^{\prime}_{4}=n_{5}\oplus n_{3}$. If $n^{\prime}_{4}+n_{0}+n_{2}\leq n_{1}+n_{3}+n_{5}$, for $n^{\prime}_{1}=n^{\prime}_{4}+n_{0}+n_{2}-n_{3}-n_{5}$, we have $n^{\prime}_{1}>n^{\prime}_{4}$ since $n_{0}+n_{2}-n_{3}-n_{5}>0$. In addition, we have $n_{3}\oplus n^{\prime}_{4}\oplus n_{5}=0,n_{0}+n_{2}+n^{\prime}_{4}=n^{\prime}_{1}+n_{3}+n_{5},n_{0}>n_{3},n_{2}>n_{5}$. Thus, $M^{\prime}=(n_{0},n^{\prime}_{1},n_{2},n_{3},n^{\prime}_{4},n_{5})$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$. On the other hand, if $n^{\prime}_{4}+n_{0}+n_{2}>n_{1}+n_{3}+n_{5}$, since $n^{\prime}_{4}\leq n_{1}$, there exist $n^{\prime}_{0},n^{\prime}_{2}$ which satisfy $n_{5}\leq n^{\prime}_{2}\leq n_{2},n_{3}\leq n^{\prime}_{0}\leq n_{0}$ and $n^{\prime}_{0}+n^{\prime}_{2}+n^{\prime}_{4}=n_{1}+n_{3}+n_{5}$. Here, $n_{3}\oplus n^{\prime}_{4}\oplus n_{5}=0,n^{\prime}_{4}\leq n_{1}$ also holds. Thus, $M^{\prime}=(n^{\prime}_{0},n_{1},n^{\prime}_{2},n_{3},n^{\prime}_{4},n_{5})$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$.
    • (i-c)
      For the case that $n_{2}\leq n_{5},n_{4}>n_{1}$, we can show $M$ has an option $M^{\prime}\in_{\circlearrowleft}P$ in a similar way to (i-b).
    • (i-d)
      Assume that $n_{2}>n_{5},n_{4}>n_{1}$. Then $n_{1}\oplus n_{3}\leq n_{5},n_{3}\oplus n_{5}\leq n_{1},$ or $n_{5}\oplus n_{1}\leq n_{3}$ holds. Without loss of generality, $n_{1}\oplus n_{3}\leq n_{5}$. Let $n^{\prime}_{2}=n_{1}\oplus n_{3}<n_{2}$. If $n_{0}+n^{\prime}_{2}+n_{4}>n_{1}+n_{3}+n_{5}$, since $n^{\prime}_{2}\leq n_{5}$, there exist $n^{\prime}_{0},n^{\prime}_{4}$ which satisfy $n_{3}\leq n^{\prime}_{0}\leq n_{0},n_{1}\leq n^{\prime}_{4}\leq n_{4}$ and $n^{\prime}_{0}+n^{\prime}_{2}+n^{\prime}_{4}=n_{1}+n_{3}+n_{5}$. Here, $n_{1}\oplus n^{\prime}_{2}\oplus n_{3}=0$ holds. Thus, $M^{\prime}=(n^{\prime}_{0},n_{1},n^{\prime}_{2},n_{3},n^{\prime}_{4},n_{5})$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$. On the other hand, if $n_{0}+n^{\prime}_{2}+n_{4}\leq n_{1}+n_{3}+n_{5}$, since $n_{0}+n_{4}>n_{1}+n_{3}$, there exists $n^{\prime}_{5}$ which satisfies $n^{\prime}_{2}<n^{\prime}_{5}\leq n_{5}$ and $n_{0}+n^{\prime}_{2}+n_{4}=n_{1}+n_{3}+n^{\prime}_{5}$. Here, $n_{1}\oplus n^{\prime}_{2}\oplus n_{3}=0,$ $n_{3}<n_{0}$, and $n_{1}<n_{4}$. Thus, $M^{\prime}=(n_{0},n_{1},n^{\prime}_{2},n_{3},n_{4},n^{\prime}_{5})$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$.
  • (ii)
    Assume that $n_{0}+n_{2}+n_{4}=n_{1}+n_{3}+n_{5}$. Then without loss of generality, we assume that $n_{0}\leq n_{3},n_{2}\leq n_{5},n_{4}\geq n_{1}$. Since $M\in N$, we have $n_{0}\oplus n_{1}\oplus n_{2}\neq 0$. Here, consider a position $(n_{0},n_{1},n_{2})$ in three-pile nim. From this position, there is a move to make a $\mathcal{P}$-position. Assume that for $n^{\prime}_{0}<n_{0}$, $(n^{\prime}_{0},n_{1},n_{2})$ is a $\mathcal{P}$-position in three-pile nim. Let $k=n_{0}-n^{\prime}_{0}$. Then, $M$ has an option $M^{\prime}=(n^{\prime}_{0},n_{1},n_{2},n_{3}-k,n_{4},n_{5})$, which satisfies $n^{\prime}_{0}+n_{2}+n_{4}=n_{1}+n_{3}-k+n_{5},n^{\prime}_{0}\oplus n_{1}\oplus n_{2}=0,n^{\prime}_{0}\leq n_{3}-k,n_{1}\leq n_{4},n_{2}\leq n_{5}$. Therefore, $M^{\prime}\in_{\circlearrowleft}P$.
Finally, it is easy to confirm that ${\rm ECN}(6_{\{1,3\}},4)$ is isomorphic to ${\rm ECN}(6_{\{1\}},4),{\rm ECN}(6_{\{2,3\}},4)$ is isomorphic to ${\rm ECN}(6_{\{2,3\}},3),{\rm ECN}(6_{\{1,2,3\}},2)$ is isomorphic to ${\rm MN}(6,2),$ and ${\rm ECN}(6_{\{1,2,3\}},4)$ is isomorphic to ${\rm ECN}(6_{\{1,2\}},4).$
From these discussions, we have characterized $\mathcal{P}$-positions in six-pile extended circular nim except for ${\rm ECN}(6_{\{1\}},2),{\rm ECN}(6_{\{1,3\}},3),{\rm ECN}(6_{\{2,3\}},2),$ and ${\rm ECN}(6_{\{1,2,3\}},3)$. Table 3 summarizes these results.
Ruleset Result
${\rm ECN}(6_{\{1\}},2)$ Unsolved
${\rm ECN}(6_{\{1\}},3)$ Shown in [6, 2]
${\rm ECN}(6_{\{1\}},4)$ Shown in [2]
${\rm ECN}(6_{\{2\}},2)$ Disjunctive sum of ${\rm MN}(3,2)$
${\rm ECN}(6_{\{2\}},i)(3\leq i\leq 4)$ Regarded as $2$-pile nim
${\rm ECN}(6_{\{3\}},i)(2\leq i\leq 4)$ Regarded as $3$-pile nim
${\rm ECN}(6_{\{1,2\}},2)$ Theorem 6
${\rm ECN}(6_{\{1,2\}},3)$ Theorem 7
${\rm ECN}(6_{\{1,2\}},4)$ Theorem 8
${\rm ECN}(6_{\{1,3\}},2)$ Theorem 9
${\rm ECN}(6_{\{1,3\}},3)$ Unsolved
${\rm ECN}(6_{\{1,3\}},4)$ Isomorphic to ${\rm ECN}(6_{\{1\}},4)$
${\rm ECN}(6_{\{2,3\}},2)$ Unsolved
${\rm ECN}(6_{\{2,3\}},3)$ Theorem 10
${\rm ECN}(6_{\{2,3\}},4)$ Isomorphic to ${\rm ECN}(6_{\{2,3\}},3)$
${\rm ECN}(6_{\{1,2,3\}},2)$ Isomorphic to ${\rm MN}(6,2)$
${\rm ECN}(6_{\{1,2,3\}},3)$ Unsolved
${\rm ECN}(6_{\{1,2,3\}},4)$ Isomorphic to ${\rm ECN}(6_{\{1,2\}},4)$
Table 3: Results of extended circular nim with six piles
2.4 Extended circular nim with seven piles
In this section, we consider extended circular nim with seven piles.
As mentioned above, circular nim with seven piles, or extended circular nim ${\rm ECN}(7_{\{1\}},i)$ is solved when $i\in\{1,4,6,7\}$.
Observation 1.
Extended circular nim ${\rm ECN}(7_{\{2\}},i)$ and ${\rm ECN}(7_{\{3\}},i)(1\leq i\leq 7)$ are isomorphic to ${\rm ECN}(7_{\{1\}},i)$.
Also, ${\rm ECN}(7_{\{2,3\}},i)$ and ${\rm ECN}(7_{\{1,3\}},i)(1\leq i\leq 7)$ are isomorphic to ${\rm ECN}(7_{\{1,2\}},i)$.
Proof.
By properly exchanging the name of the piles, we can easily show the isomorphism of rulesets.
Figure 3 shows the case of ${\rm ECN}(7_{\{1,2\}},i)$ and ${\rm ECN}(7_{\{1,3\}},i)$. ∎
Figure 3: ${\rm ECN}(7_{\{1,2\}},i)$ and ${\rm ECN}(7_{\{1,3\}},i)$ are isomorphic
Therefore, without loss of generality, we only consider ${\rm ECN}(7_{\{1,2\}},i),$ and ${\rm ECN}(7_{\{1,2,3\}},i)(2\leq i\leq 5)$.
Theorem 11.
Consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6})$ in ${\rm ECN}(7_{\{1,2\}},4)$. Then, $M$ is a $\mathcal{P}$-position if and only if $M\in_{\circlearrowleft}P=\{(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6})\mid n_{0}=n_{1}=n_{4},n_{2}=n_{6},n_{3}+n_{5}=n_{0}+n_{2},n_{0}=\min(\{n_{i}\})\}$.
Proof.
Let $N=\{M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6})\mid M\not\in_{\circlearrowleft}P\}$ The terminal position is in $P$.
Next, consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6})=(a,a,a+b,a+c,a,a+d,a+b)(c+d=b)\in P$. The player cannot remove all tokens from $n_{2},n_{3},n_{5},n_{6}$. Let $M^{\prime}=(n^{\prime}_{0},n^{\prime}_{1},n^{\prime}_{2},n^{\prime}_{3},n^{\prime}_{4},n^{\prime}_{5},n^{\prime}_{6})$ be an option of $M$. Assume that $\min(\{n^{\prime}_{i}\})=n_{0}=a$. For $M^{\prime}=(a,a,a+b^{\prime},a+c^{\prime},a,a+d^{\prime},a+b)(b^{\prime}+c^{\prime}+d^{\prime}<b+c+d,b^{\prime}\leq b,c^{\prime}\leq c,d^{\prime}\leq d)$, we can say that $M^{\prime}\not\in_{\circlearrowleft}P$ because if $b=b^{\prime}$ then $c^{\prime}+d^{\prime}<b$, if $b=c^{\prime}$ then $c^{\prime}=c=b,d=d^{\prime}=0,b^{\prime}+0=b^{\prime}<b,$ and if $b^{\prime}=d^{\prime}$ then $b+0<b^{\prime}$ or $b+0=b^{\prime}=d^{\prime},c=c^{\prime}=0$. For $M^{\prime}=(a,a,a+b,a+c^{\prime},a,a+d^{\prime},a+b^{\prime})(b^{\prime}+c^{\prime}+d^{\prime}<b+c+d,b^{\prime}\leq b,c^{\prime}\leq c,d^{\prime}\leq d)$, we can say that $M^{\prime}\not\in_{\circlearrowleft}P$ in a similar way. Also, for $M^{\prime}=(a,a,a+b^{\prime},a+c,a+d,a+b^{\prime\prime})(b^{\prime}+b^{\prime\prime}<2b,b^{\prime},b^{\prime\prime}\leq b),$ we can say that $M^{\prime}\not\in_{\circlearrowleft}P$ because if $b^{\prime}=b^{\prime\prime}$ then $c+d=b>b^{\prime},$ if $b^{\prime}=d$ then $c>0$ or $b^{\prime\prime}<b$, and if $b^{\prime\prime}=c$ then $d>0$ or $b^{\prime}<b$.
Assume that $\min(\{n^{\prime}_{i}\})<n_{0}$. If $M^{\prime}\in_{\circlearrowleft}P$, then there exists $i$ such that $n^{\prime}_{i}=n^{\prime}_{((i+1)\bmod 7)}=n^{\prime}_{((i+4)\bmod 7)}<n_{0}$, but selecting these three piles is not legal in ${\rm ECN}(7_{1,2},4)$. Therefore, $M^{\prime}\in N$.
Also, for any position $M\not\in P$ but $M\in_{\circlearrowleft}P$, we can show that every option $M^{\prime}\in N$ in a similar way.
Finally, consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6})\in N$. We assume $n_{0}=\min(\{n_{i}\})$. Then $M=(a,a+a_{1},a+a_{2},a+a_{3},a+a_{4},a+a_{5},a+a_{6})$.
Consider the case of $a_{2}\geq a_{5}$.
  • (i)
    Assume that $a_{2}\geq a_{5},a_{6}>a_{5}$.
    • (i-a)
      If $a_{2}\geq a_{6},a_{3}+a_{5}\geq a_{6}$, then $M^{\prime}=(a,a,a+a_{6},a+a_{6}-a_{5},a,a+a_{5},a+a_{6})$ is an option of $M$ and $M^{\prime}\in P$.
    • (i-b)
      If $a_{2}\geq a_{6},a_{3}+a_{5}<a_{6}$, then $M^{\prime}=(a,a,a+a_{3}+a_{5},a+a_{3},a,a+a_{5},a+a_{3}+a_{5})$ is an option of $M$ and $M^{\prime}\in P$.
    • (i-c)
      If $a_{2}<a_{6},a_{3}+a_{5}\geq a_{2}$, then $M^{\prime}=(a,a,a+a_{2},a+a_{2}-a_{5},a,a+a_{5},a+a_{2})$ is an option of $M$ and $M^{\prime}\in P$.
    • (i-d)
      If $a_{2}<a_{6},a_{3}+a_{5}<a_{2}$, then $(a,a,a+a_{3}+a_{5},a+a_{3},a,a+a_{5},a+a_{3}+a_{5})$ is an option of $M$ and $M^{\prime}\in P$.
  • (ii)
    If $a_{2}\geq a_{5},a_{6}\leq a_{5},a_{1}+a_{6}\geq a_{5}$, then $M^{\prime}=(a,a+a_{5}-a_{6},a+a_{5},a,a,a+a_{5},a+a_{6})$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$.
  • (iii)
    If $a_{2}\geq a_{5},a_{6}\leq a_{5},a_{1}+a_{6}<a_{5}$, then $M^{\prime}=(a,a+a_{1},a+a_{1}+a_{6},a,a,a+a_{1}+a_{6},a+a_{6})$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$.
For the case of $a_{5}<a_{2}$, we can have a similar discussion. Therefore, $M\in N$ has an option $M^{\prime}\in_{\circlearrowleft}P$.
Theorem 12.
Consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6})$ in ${\rm ECN}(7_{\{1,2\}},5)$. $M$ is a $\mathcal{P}$-position if and only if $M\in_{\circlearrowleft}P=\{(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6})\mid n_{0}=0,n_{1}=n_{2}=n_{3}+n_{4}=n_{5}=n_{6}\}$.
Proof.
Let $N=\{M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6})\mid M\not\in_{\circlearrowleft}P\}$. The terminal position is in $P$.
Next, consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6})\in P$. From the definition of $P$, for any position $M\in P$, $5\times\max(\{n_{i}\})=\sum{n_{i}}$ holds.
Let $M^{\prime}=(n^{\prime}_{0},n^{\prime}_{1},n^{\prime}_{2},n^{\prime}_{3},n^{\prime}_{4},n^{\prime}_{5},n^{\prime}_{6})$ be an option of $M$. If $n^{\prime}_{1}=n_{1},n^{\prime}_{2}=n_{2},n^{\prime}_{5}=n_{5},$ or $n^{\prime}_{6}=n_{6},$ then $\max(\{n_{i}\})=\max(\{n^{\prime}_{i}\})$, but $\sum{n^{\prime}_{i}}<\sum{n_{i}}$. Therefore, $5\times\max(\{n^{\prime}_{i}\})>\sum{n^{\prime}_{i}}$ and $M^{\prime}\not\in P$.
If $n^{\prime}_{1}<n_{1},n^{\prime}_{2}<n_{2},n^{\prime}_{5}<n_{5},$ and $n^{\prime}_{6}<n_{6},$ then $n^{\prime}_{3}=n_{3}$ and $n^{\prime}_{4}=n_{4}$.
Since $n^{\prime}_{3}+n^{\prime}_{4}=n_{3}+n_{4}=n_{1}>n^{\prime}_{1}$, we do not have $n^{\prime}_{0}=0,n^{\prime}_{1}=n^{\prime}_{2}=n^{\prime}_{3}+n^{\prime}_{4}=n^{\prime}_{5}=n^{\prime}_{6}$.
Also, we do not have $n^{\prime}_{1}=0,n^{\prime}_{2}=n^{\prime}_{3}=n^{\prime}_{4}+n^{\prime}_{5}=n^{\prime}_{6}=n^{\prime}_{0}$ or $n^{\prime}_{2}=0,n^{\prime}_{3}=n^{\prime}_{4}=n^{\prime}_{5}+n^{\prime}_{6}=n^{\prime}_{0}=n^{\prime}_{1}$ or $n^{\prime}_{5}=0,n^{\prime}_{6}=n^{\prime}_{0}=n^{\prime}_{1}+n^{\prime}_{2}=n^{\prime}_{3}=n^{\prime}_{4}$ or $n^{\prime}_{6}=0,n^{\prime}_{0}=n^{\prime}_{1}=n^{\prime}_{2}+n^{\prime}_{3}=n^{\prime}_{4}=n^{\prime}_{5}$ because they contradict $n^{\prime}_{0}=n_{0}=0$ and $n^{\prime}_{3}+n^{\prime}_{4}>0$.
The case $n^{\prime}_{3}=0,n^{\prime}_{4}=n^{\prime}_{5}=n^{\prime}_{6}+n^{\prime}_{0}=n^{\prime}_{1}=n^{\prime}_{2}$ also contradicts because $n^{\prime}_{3}+n^{\prime}_{4}=n_{3}+n_{4}=n_{5}$ and $n^{\prime}_{4}=n_{5}>n^{\prime}_{5}$. In a similar way, $n^{\prime}_{4}=0,n^{\prime}_{5}=n^{\prime}_{6}=n^{\prime}_{0}+n^{\prime}_{1}=n^{\prime}_{2}=n^{\prime}_{3}$ contradicts, too.
Therefore, for $M\in P,$ every option $M^{\prime}\in N$.
Also, for any position $M\not\in P$ but $M\in_{\circlearrowleft}P$, we can show that every option $M^{\prime}\in N$ in a similar way.
Finally, consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6})\not\in N$.
Without loss of generality, we assume that $\min(\{n_{i}\})=n_{0}$. We consider three cases:
  • (i)
    Assume that $\min(\{n_{1},n_{2},n_{3},n_{4},n_{5},n_{6}\})=n_{1}$. $M^{\prime}=(n_{0},n_{1},n_{1},0,n_{1},n_{1},n_{1}-n_{0})$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$. For the case of $\min(\{n_{1},n_{2},n_{3},n_{4},n_{5},n_{6}\})=n_{6}$, we have an option $M^{\prime}\in_{\circlearrowleft}P$ in a similar way.
  • (ii)
    Assume that $\min(\{n_{1},n_{2},n_{3},n_{4},n_{5},n_{6}\})=n_{2}$. $M^{\prime}=(n_{0},n_{2}-n_{0},n_{2},n_{2},0,n_{2},n_{2})$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$. For the case of $\min(\{n_{1},n_{2},n_{3},n_{4},n_{5},n_{6}\})=n_{5}$, we have an option $M^{\prime}\in_{\circlearrowleft}P$ in a similar way.
  • (iii)
    Assume that $\min(\{n_{1},n_{2},n_{3},n_{4},n_{5},n_{6}\})=n_{3}$.
    • (iii-1)
      If $\min(\{n_{1},n_{2},n_{4},n_{5},n_{6}\})=n_{i}(i\in\{1,2,5\})$, $M^{\prime}=(0,n_{i},n_{i},n_{3},n_{i}-n_{3},n_{i},n_{i})$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$.
    • (iii-2)
      If $\min(\{n_{1},n_{2},n_{4},n_{5},n_{6}\})=n_{6}$ and $n_{0}+n_{6}\leq\min(\{n_{1},n_{2},n_{4},n_{5}\})$, $M^{\prime}=(n_{0},n_{0}+n_{6},n_{0}+n_{6},0,n_{0}+n_{6},n_{0}+n_{6},n_{6})$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$.
    • (iii-3)
      Consider the case that $\min(\{n_{1},n_{2},n_{4},n_{5},n_{6}\})=n_{6}$ and $n_{0}+n_{6}>\min(\{n_{1},n_{2},n_{4},n_{5}\})$. Let $a=\min(\{n_{1},n_{2},n_{4},n_{5}\})$. $n_{0}\leq a,n_{6}\leq a,n_{0}+n_{6}>a$ hold.
      • (iii-3-a)
        If $n_{1}=a$, $n_{2}=a$, or $n_{5}=a$, $M^{\prime}=(n_{0},a,a,0,a,a,a-n_{0})$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$.
      • (iii-3-b)
        If $n_{4}=a$, $M^{\prime}=(a-n_{6},a,a,0,a,a,n_{6})$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$.
    • (iii-4)
      If $\min(\{n_{1},n_{2},n_{4},n_{5},n_{6}\})=n_{4}$ and $n_{3}+n_{4}\leq\min(\{n_{1},n_{2},n_{5},n_{6}\})$, $M^{\prime}=(0,n_{3}+n_{4},n_{3}+n_{4},n_{3},n_{4},n_{3}+n_{4},n_{3}+n_{4})$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$.
    • (iii-5)
      Consider the case that $\min(\{n_{1},n_{2},n_{4},n_{5},n_{6}\})=n_{4}$ and $n_{3}+n_{4}>\min(\{n_{1},n_{2},n_{5},n_{6}\})$. Let $a=\min(\{n_{1},n_{2},n_{5},n_{6}\})$. $n_{3}\leq a,n_{4}\leq a,n_{3}+n_{4}>a$ hold.
      • (iii-5-a)
        If $n_{2}=a$, $n_{5}=a$, or $n_{6}=a$, then $M^{\prime}=(0,a,a,a-n_{4},n_{4},a,a)$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$.
      • (iii-5-b)
        If $n_{1}=a$, $M^{\prime}=(0,a,a,n_{3},a-n_{3},a,a)$ is an option of $M$ and $M^{\prime}\in_{\circlearrowleft}P$.
    For the case of $\min(\{n_{1},n_{2},n_{3},n_{4},n_{5},n_{6}\})=n_{4}$, we have an option $M^{\prime}\in_{\circlearrowleft}P$ in a similar way.
Observation 2.
Extended circular nim ${\rm ECN}(7_{\{1,2,3\}},2)$ is isomorphic to Moore’s nim ${\rm MN}(7,2)$. Also, Extended circular nim ${\rm ECN}(7_{\{1,2,3\}},2)$ is isomorphic to ${\rm MN}(7,5)$.
Proof.
In ${\rm ECN}(7_{\{1,2,3\}},2)$, the player can choose any $2$ piles. Also, in ${\rm ECN}(7_{\{1,2,3\}},5)$, the player can choose any $5$ piles. ∎
From these discussions, we have characterized $\mathcal{P}$-positions in seven-pile extended circular nim except for ${\rm ECN}(7_{\{1\}},2),{\rm ECN}(7_{\{1\}},3),{\rm ECN}(7_{\{1\}},5),{\rm ECN}(7_{\{1,2\}},2),\allowbreak{\rm ECN}(7_{\{1,2\}},3),{\rm ECN}(7_{\{1,2,3\}},3),{\rm ECN}(7_{\{1,2,3\}},4)$, and those which are isomorphic to them. Table 4 summarizes these results.
Ruleset Result
${\rm ECN}(7_{\{1\}},i)(2\leq i\leq 3)$ Unsolved
${\rm ECN}(7_{\{1\}},4)$ Shown in [3]
${\rm ECN}(7_{\{1\}},5)$ Unsolved
${\rm ECN}(7_{\{2\}},i)(2\leq i\leq 5)$ Isomorphic to ${\rm ECN}(7_{\{1\}},i)$
${\rm ECN}(7_{\{3\}},i)(2\leq i\leq 5)$ Isomorphic to ${\rm ECN}(7_{\{1\}},i)$
${\rm ECN}(7_{\{1,2\}},i)(2\leq i\leq 3)$ Unsolved
${\rm ECN}(7_{\{1,2\}},4)$ Theorem 11
${\rm ECN}(7_{\{1,2\}},5)$ Theorem 12
${\rm ECN}(7_{\{1,3\}},i)(2\leq i\leq 5)$ Isomorphic to ${\rm ECN}(7_{\{1,2\}},i)$
[…]
${\rm ECN}(7_{\{1,2,3\}},2)$ Isomorphic to ${\rm MN}(7,2)$
${\rm ECN}(7_{\{1,2,3\}},i)(3\leq i\leq 4)$ Unsolved
${\rm ECN}(7_{\{1,2,3\}},5)$ Isomorphic to ${\rm MN}(7,5)$
Table 4: Results of extended circular nim with seven piles
2.5 Extended circular nim with eight piles
For extended circular nim with eight piles, many cases are still open problems, but we can characterize $\mathcal{P}$-positions for some cases.
We consider ${\rm ECN}(8_{S},i)(2\leq i\leq 6)$. As mentioned above, circular nim with eight piles, or extended circular nim ${\rm ECN}(8_{\{1\}},i)$ is solved when $i\in\{1,6,7,8\}$.
When $S=\{2\},$ the position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6},n_{7},n_{8})$ can be considered as a disjunctive sum of positions $(n_{0},n_{2},n_{4},n_{6})$ and $(n_{1},n_{3},n_{5},n_{7})$ in circular nim ${\rm CN}(4,i)$. If $i=2$, $M$ is disjunctive sum of two positions in ${\mathrm{C}N}(4,2)$, but any closed formula for Sprague-Grundy values of ${\mathrm{C}N}(4,2)$ is unknown. Therefore, it is an open problem to characterize $\mathcal{P}$-positions in ${\rm ECN}(8_{\{2\}},2)$. If $i=3,$ $M$ is disjunctive sum of two positions in ${\mathrm{C}N}(4,3)$. The formula for Sprague-Grundy value is shown in [7], thus this case is solved. If $i\geq 4,$ $M$ is the disjunctive sum of positions in ${\mathrm{C}N}(4,i)$. In this case, each player can choose all four piles which have $n_{0},n_{2},n_{4},n_{6}$ tokens or all four piles which have $n_{1},n_{3},n_{5},n_{7}$ tokens. Therefore, this position can be considered as a position $(n_{0}+n_{2}+n_{4}+n_{6},n_{1}+n_{3}+n_{5}+n_{7})$ in two-pile nim and $M$ is a $\mathcal{P}$-position if and only if $n_{0}+n_{2}+n_{4}+n_{6}=n_{1}+n_{3}+n_{5}+n_{7}$.
When $S=\{3\}$ ${\rm ECN}(8_{3},i)$ is isomorphic to ${\rm ECN}(8_{1},i)$. As shown in Figure 4, we have this isomorphism by exchanging names of piles.
Figure 4: ${\rm ECN}(8_{\{1\}},i)$ and ${\rm ECN}(8_{\{3\}},i)$ are isomorphic
When $S=\{4\}$ and $i\geq 2$, ${\rm ECN}(8_{\{4\}},i)$ can be considered as a position $(n_{0}+n_{4},n_{1}+n_{5},n_{2}+n_{6},n_{3}+n_{7})$ in four-pile nim and $M$ is a $\mathcal{P}$-position if and only if $(n_{0}+n_{4})\oplus(n_{1}+n_{5})\oplus(n_{2}+n_{6})\oplus(n_{3}+n_{7})=0$
Theorem 13.
Consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6},n_{7})$ in ${\rm ECN}(8_{\{1,3\}},2)$. $M$ is a $\mathcal{P}$-position if and only if $M\in S=\{(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6},n_{7})\mid n_{0}\oplus n_{2}\oplus n_{4}\oplus n_{6}=n_{1}\oplus n_{3}\oplus n_{5}\oplus n_{7}=0\}$.
Proof.
In ${\rm ECN}(8_{\{1,3\}},2)$, the player can reduce at most one of $n_{0},n_{2},n_{4},n_{6}$, and at most one of $n_{1},n_{3},n_{5},n_{7}$ in one move. Therefore, this ruleset is selective sum of two positions $(n_{0},n_{2},n_{4},n_{6}),$ and $(n_{1},n_{3},n_{5},n_{7})$ in four-pile nim. Thus, from Theorem 4, $M$ is a $\mathcal{P}$-position if and only if such two four-pile nim positions are $\mathcal{P}$-positions. ∎
Theorem 14.
Consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6},n_{7})$ in ${\rm ECN}(8_{\{1,3\}},4)$. $M$ is a $\mathcal{P}$-position if and only if $M\in P=\{(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6},n_{7})\mid n_{0}=n_{4},n_{1}=n_{5},n_{2}=n_{6},n_{3}=n_{7}\}$.
Proof.
In ${\rm ECN}(8_{\{1,3\}},4)$, the player can reduce at most one of $n_{0},n_{4}$, at most one of $n_{1},n_{5}$, at most one of $n_{2},n_{6},$ and at most one of $n_{3},n_{7}$ in one move. Therefore, this ruleset is the selective sum of four positions $(n_{0},n_{4}),(n_{1},n_{5}),(n_{2},n_{6}),$ and $(n_{3},n_{7})$ in two-pile nim. Thus, from Theorem 4, $M$ is a $\mathcal{P}$-position if and only if such four two-pile nim positions are $\mathcal{P}$-positions. ∎
Theorem 15.
Consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6},n_{7})$ in ${\rm ECN}(8_{\{1,3\}},6)$. $M$ is a $\mathcal{P}$-position if and only if $M\in P=\{(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6},n_{7})\mid n_{0}=n_{2}=n_{4}=n_{6},n_{1}=n_{3}=n_{5}=n_{7}\}$.
Proof.
In ${\rm ECN}(8_{\{1,3\}},6)$, the player can reduce at most three of $n_{0},n_{2},n_{4},n_{6}$ and at most three of $n_{1},n_{3},n_{5},n_{7}$ in one move. Therefore, this ruleset is the selective sum of two positions $(n_{0},n_{2},n_{4},n_{6})$ and $(n_{1},n_{3},n_{5},n_{7})$ in ${\rm MN}(4,3)$. Thus, from Theorem 4, $M$ is a $\mathcal{P}$-position if and only if such two Moore’s nim positions are $\mathcal{P}$-positions. ∎
When $S=\{1,4\}$ and $i\leq 4$, ${\rm ECN}(8_{\{1,4\}},i)$ has not been solved. For the case of $i\geq 5$, ${\rm ECN}(8_{\{1,4\}},i)$ is isomorphic to ${\rm ECN}(8_{\{1\}},i)$.
When $S=\{2,3\},$ similar to the case of $S=\{3\}$, by exchanging the names of the piles, we can show that ${\rm ECN}(8_{\{2,3\}},i)$ is isomorphic to ${\rm ECN}(8_{\{1,2\}},i)$.
When $S=\{2,4\}$, similar to the case of $S=\{2\}$, the position is a disjunctive sum of two positions $(n_{0},n_{2},n_{4},n_{6})$ and $(n_{1},n_{3},n_{5},n_{7})$ in ${\rm ECN}(4_{\{1,2\}},\min(i,4))$.
When $S=\{3,4\}$, similar to the case of $S=\{3\}$, by exchanging the names of the piles, we can show that ${\rm ECN}(8_{\{3,4\}},i)$ is isomorphic to ${\rm ECN}(8_{\{1,4\}},i)$.
When $S=\{1,2,3\}$ and $2\leq i\leq 5$, ${\rm ECN}(8_{\{1,2,3\}},i)$ has not been solved. For the case of $i=6$, we have the following theorem.
Theorem 16.
Consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6},n_{7})$ in ${\rm ECN}(8_{\{1,2,3\}},6)$. $M$ is a $\mathcal{P}$-position if and only if $M\in_{\circlearrowleft}P=\{(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6},n_{7})\mid n_{0}=n_{2}=n_{4}=n_{6}=n_{1}+n_{3}+n_{5}+n_{7}\}\setminus\{(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6},n_{7})\mid n_{0}=n_{2}=n_{4}=n_{6}\text{ and }n_{1}=n_{3}=n_{5}=n_{7}\}\cup\{(0,0,0,0,0,0,0,0)\}$
Proof.
Let $N=\{M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6})\mid M\not\in_{\circlearrowleft}P\}$. The terminal position is in $P$.
Consider a non-terminal position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6},n_{7})\in P$.
There are three types of moves:
  • (i)
    Consider the case that the player selects $n_{0},n_{2},n_{4},n_{6}$ or part of them to reduce. Let $M^{\prime}=(n^{\prime}_{0},n_{1},n^{\prime}_{2},n_{3},n^{\prime}_{4},n_{5},n^{\prime}_{6},n_{7})(n^{\prime}_{0}+n^{\prime}_{2}+n^{\prime}_{4}+n^{\prime}_{6}<n_{0}+n_{2}+n_{4}+n_{6})$ is an option of $M$. If $n^{\prime}_{0}=n^{\prime}_{2}=n^{\prime}_{4}=n^{\prime}_{6}$, then $n_{1}+n_{3}+n_{5}+n_{7}>n^{\prime}_{0}$. Also, since $M\in P$ and $n_{0}=n_{2}=n_{4}=n_{6}$, we do not have $n_{1}=n_{3}=n_{5}=n_{7}$. Therefore, $M^{\prime}\in N$.
  • (ii)
    Consider the case that the player selects $n_{1},n_{3},n_{5},n_{7}$ or part of them to reduce. Let $M^{\prime}=(n_{0},n^{\prime}_{1},n_{2},n^{\prime}_{3},n_{4},n^{\prime}_{5},n_{6},n^{\prime}_{7})(n^{\prime}_{1}+n^{\prime}_{3}+n^{\prime}_{5}+n^{\prime}_{7}<n_{1}+n_{3}+n_{5}+n_{7})$ is an option of $M$. Since $n_{0}=n_{2}=n_{4}=n_{6}>n^{\prime}_{1}+n^{\prime}_{3}+n^{\prime}_{5}+n^{\prime}_{7}$, $M\in N$.
  • (iii)
    Consider the case that the player selects at least one and at most three of $n_{0},n_{2},n_{4},n_{6}$ and at least one and at most three of $n_{1},n_{3},n_{5},n_{7}$ to reduce.
    Let $M^{\prime}=(n^{\prime}_{0},n^{\prime}_{1},n^{\prime}_{2},n^{\prime}_{3},n^{\prime}_{4},n^{\prime}_{5},n^{\prime}_{6},n^{\prime}_{7})(n^{\prime}_{0}+n^{\prime}_{2}+n^{\prime}_{4}+n^{\prime}_{6}<n_{0}+n_{2}+n_{4}+n_{6},n^{\prime}_{1}+n^{\prime}_{3}+n^{\prime}_{5}+n^{\prime}_{7}<n_{1}+n_{3}+n_{5}+n_{7})$ is an option of $M$ after such a move. Then, we do not have $n^{\prime}_{0}=n^{\prime}_{2}=n^{\prime}_{4}=n^{\prime}_{6}$. Also, since $\max(n^{\prime}_{0},n^{\prime}_{2},n^{\prime}_{4},n^{\prime}_{6})=n_{0}=n_{1}+n_{3}+n_{5}+n_{7}>n^{\prime}_{1}+n^{\prime}_{3}+n^{\prime}_{5}+n^{\prime}_{7}$, we do not have $n^{\prime}_{1}=n^{\prime}_{3}=n^{\prime}_{5}=n^{\prime}_{7}=n^{\prime}_{0}+n^{\prime}_{2}+n^{\prime}_{4}+n^{\prime}_{6}$. Therefore, $M\in N$.
Also, for any position $M\not\in P$ but $M\in_{\circlearrowleft}P$, we can show that every option $M^{\prime}\in N$ in a similar way.
Finally, consider a position $M=(n_{0},n_{1},n_{2},n_{3},n_{4},n_{5},n_{6},n_{7})\in N$.
If $n_{0}=n_{2}=n_{4}=n_{6}=n_{1}+n_{3}+n_{5}+n_{7},n_{1}=n_{3}=n_{5}=n_{7}$, then $M^{\prime}=(n_{1},n_{1},0,n_{3},0,n_{5},0,n_{7})$ is an option of $M$ and $M^{\prime}\in P$.
Otherwise, without loss of generality, we assume $\min(\{n_{0},n_{2},n_{4},n_{6}\})<\min(\{n_{1},n_{3},n_{5},n_{7}\})$.
  • (i)
    Consider the case that $n_{0}+n_{2}+n_{4}+n_{6}\leq\min(\{n_{1},n_{3},n_{5},n_{7}\})$. If we have $n_{0}=n_{2}=n_{4}=n_{6}$, $M^{\prime}=(n_{0},n_{0},n_{2},0,n_{4},0,n_{6},0)$ is an option of $M$ and $M^{\prime}\in P$. If we do not have $n_{0}=n_{2}=n_{4}=n_{6}$, $M^{\prime}=(n_{0},n_{0}+n_{2}+n_{4}+n_{6},n_{2},n_{0}+n_{2}+n_{4}+n_{6},n_{4},n_{0}+n_{2}+n_{4}+n_{6},n_{6},n_{0}+n_{2}+n_{4}+n_{6})$ is an option of $M$ and $M^{\prime}\in P$.
  • (ii)
    Consider the case that $n_{0}+n_{2}+n_{4}+n_{6}>\min(\{n_{1},n_{3},n_{5},n_{7}\})$. Without loss of generality, $n_{1}=\min(\{n_{1},n_{3},n_{5},n_{7}\})$. For the case of $n_{0}=\min(\{n_{0},n_{2},n_{4},n_{6}\}),$ since $n_{1}>n_{0}=\min(\{n_{0},n_{2},n_{4},n_{6}\})$, we have $(n_{0},n^{\prime}_{2},n^{\prime}_{4},n^{\prime}_{6})$ such that $n_{0}+n^{\prime}_{2}+n^{\prime}_{4}+n^{\prime}_{6}=n_{1},n^{\prime}_{2}\leq n_{2},n^{\prime}_{4}\leq n_{4},n^{\prime}_{6}\leq n_{6},$ and not satisfy $n_{0}=n^{\prime}_{2}=n^{\prime}_{4}=n^{\prime}_{6}$. Therefore, $M^{\prime}=(n_{0},n_{1},n^{\prime}_{2},n_{1},n^{\prime}_{4},n_{1},n^{\prime}_{6},n_{1})$ is an option of $M$ and $M^{\prime}\in P$.
    Similarly, for the cases of $n_{2}=\min(\{n_{0},n_{2},n_{4},n_{6}\}),n_{4}=\min(\{n_{0},n_{2},n_{4},n_{6}\}),$ and $n_{6}=\min(\{n_{0},n_{2},n_{4},n_{6}\}),$ we have an option $M^{\prime}\in P$.
When $S=\{1,2,4\}$, ${\rm ECN}(8_{\{1,2,4\}},2)$ has not been solved. For the case of $3\leq i\leq 6$, ${\rm ECN}(8_{\{1,2,4\}},i)$ is isomorphic to ${\rm ECN}(8_{\{1,2\}},i)$.
When $S=\{1,3,4\}$ and $2\leq i\leq 4$, ${\rm ECN}(8_{\{1,3,4\}},i)$ has not been solved. For the case of $5\leq i\leq 6$, ${\rm ECN}(8_{\{1,3,4\}},i)$ is isomorphic to ${\rm ECN}(8_{\{1,3\}},i)$.
When $S=\{2,3,4\}$, similar to the case of $S=\{3\}$, by exchanging the names of the piles, we can show that ${\rm ECN}(8_{\{2,3,4\}},i)$ is isomorphic to ${\rm ECN}(8_{\{1,2,4\}},i)$.
When $S=\{1,2,3,4\}$, if $i=2$, ${\rm ECN}(8_{\{1,2,3,4\}},2)$ is isomorphic to ${\rm MN}(8,2)$. For the case of $3\leq i\leq 6$, ${\rm ECN}(8_{\{1,2,3,4\}},2)$ is isomorphic to ${\rm ECN}(8_{\{1,2,3\}},i)$.
Table 5 summarizes results shown in this section.
Ruleset Result
${\rm ECN}(8_{\{1\}},i)(2\leq i\leq 5)$ Unsolved
${\rm ECN}(8_{\{1\}},6)$ Shown in [2]
${\rm ECN}(8_{\{2\}},i)(2\leq i\leq 6)$ Disjunctive sum of ${\rm CN}(4,\min(\{4,i\}))$
${\rm ECN}(8_{\{3\}},i)(2\leq i\leq 6)$ Isomorphic to ${\rm ECN}(8_{\{1\}},i)$
${\rm ECN}(8_{\{4\}},i)(2\leq i\leq 6)$ Regarded as $4$-pile nim
${\rm ECN}(8_{\{1,2\}},i)(2\leq i\leq 6)$ Unsolved
${\rm ECN}(8_{\{1,3\}},2)$ Theorem 13
${\rm ECN}(8_{\{1,3\}},3)$ Unsolved
${\rm ECN}(8_{\{1,3\}},4)$ Theorem 14
${\rm ECN}(8_{\{1,3\}},5)$ Unsolved
${\rm ECN}(8_{\{1,3\}},6)$ Theorem 15
${\rm ECN}(8_{\{1,4\}},i)(2\leq i\leq 4)$ Unsolved
${\rm ECN}(8_{\{1,4\}},i)(5\leq i\leq 6)$ Isomorphic to ${\rm ECN}(8_{\{1\}},i)$
${\rm ECN}(8_{\{2,3\}},i)(2\leq i\leq 6)$ Isomorphic to ${\rm ECN}(8_{\{1,2\}},i)$
${\rm ECN}(8_{\{2,4\}},i)(2\leq i\leq 6)$ Disjunctive sum of ${\rm ECN}(4_{\{1,2\}},\min(\{4,i\}))$
${\rm ECN}(8_{\{3,4\}},i)(2\leq i\leq 6)$ Isomorphic to ${\rm ECN}(8_{\{1,4\}},i)$
${\rm ECN}(8_{\{1,2,3\}},i)(2\leq i\leq 5)$ Unsolved
${\rm ECN}(8_{\{1,2,3\}},6)$ Theorem 16
${\rm ECN}(8_{\{1,2,4\}},2)$ Unsolved
${\rm ECN}(8_{\{1,2,4\}},i)(3\leq i\leq 6)$ Isomorphic to ${\rm ECN}(8_{\{1,2\}},i)$
${\rm ECN}(8_{\{1,3,4\}},i)(2\leq i\leq 4)$ Unsolved
${\rm ECN}(8_{\{1,3,4\}},i)(5\leq i\leq 6)$ Isomorphic to ${\rm ECN}(8_{\{1,3\}},i)$
${\rm ECN}(8_{\{2,3,4\}},i)(2\leq i\leq 6)$ Isomorphic to ${\rm ECN}(8_{\{1,2,3\}},i)$
${\rm ECN}(8_{\{1,2,3,4\}},2)$ Isomorphic to ${\rm MN}(8,2)$
${\rm ECN}(8_{\{1,2,3,4\}},i)(3\leq i\leq 6)$ Isomorphic to ${\rm ECN}(8_{\{1,2,3\}},i)$
Table 5: Results of extended circular nim with eight piles
2.6 Further generalization
From the results on ${\rm ECN}(5_{1},3)$ and ${\rm ECN}(7_{1,2},5)$, we have a generalization as follows:
Theorem 17.
Let $m(>1)$ be an integer and assume that $12m+$ 1 is a prime number. Consider a position $M=(n_{0},n_{1},\ldots,n_{2m})$ in ${\rm ECN}((2m+1)_{\{1,2,\ldots,m-1\}},2m-1)$. $M$ is a $\mathcal{P}$-position if and only if $M\in_{\circlearrowleft}P=\{M\mid n_{0}=0,n_{1}=n_{2}=\cdots=n_{m-1}=n_{m}+n_{m+1}=n_{m+2}=\cdots=n_{2m}\}$.
Proof.
The proof is similar to the proof of Theorem 12. ∎
Also, we have a generalization from the results on ${\rm ECN}(4_{\{1\}},2),{\rm ECN}(6_{\{1,3\}},2),$ and ${\rm ECN}(8_{\{1,3\}},2)$, and a generalization from the results on ${\rm ECN}(4_{\{1\}},2)$ and ${\rm ECN}(8_{\{1,3\}},6)$.
Theorem 18.
Let $m(>1)$ be an integer and assume that $h$ is maximal odd integer not more than $m$. Consider a position $M=(n_{0},n_{1},\ldots,n_{2m-1})$ in ${\rm ECN}((2m)_{\{1,3,\ldots,h\}},2)$. Then, $M$ is a $\mathcal{P}$-position if and only if $M\in P=\{(n_{0},n_{1},\ldots,n_{2m-1})\mid n_{0}\oplus n_{2}\oplus\cdots\oplus n_{2m-2}=n_{1}\oplus n_{3}\oplus\cdots\oplus n_{2m-1}=0\}$.
Proof.
The proof is similar to the proof of Theorem 13. ∎
Theorem 19.
Let $m(>1)$ be an integer and assume that $h=2^{m-1}-1$. Consider a position $M=(n_{0},n_{1},\ldots,n_{2^{m}-1})$ in ${\rm ECN}((2^{m})_{\{1,3,\ldots,h\}},2^{m}-2)$. Then, $M$ is a $\mathcal{P}$-position if and only if $M\in P=\{(n_{0},n_{1},\ldots,n_{2^{m}-1})\mid n_{0}=n_{2}=\cdots=n_{2^{m}-2},n_{1}=n_{3}=\cdots=n_{2^{m}-1}\}$.
Proof.
The proof is similar to the proof of Theorem 15. ∎